OpenStudy (anonymous):

what is the solution to the problem : 4 over 5 plus negative 3 over 5?

OpenStudy (anonymous):

-7 OVER 5

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

did i get star

OpenStudy (anonymous):

Shouldnt it be \[\frac{ 4 }{5 }+(-\frac{ 3 }{ 5 })\] \[\frac{ 4-3 }{ 5 }=\frac{ 1 }{5 }\]

OpenStudy (shamil98):

yeah, i'm being dumb today, -_- sorry